Sara wants to have $ 580,000 in her savings account when she retires. How much must she put in the account​ now, if the account pays a fixed interest rate of 10​%, to ensure that she has $ 580,000 in 18 years​ time?

A. $187,772

B. $146,045

C. $110,817

D. $104,318
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

580,000=P(1.1)^18

P=580,000/(1.1)^18

=$104318(Approx).